Impulse and momentum
- a 100kg car is moving w/ an initial velocity of 22m/s in a time interval of 1.5sec before coming to stop. What is the force of the moving car?
1
Expert's answer
2018-10-22T11:22:09-0400

We can find the force on the moving car from the definition of the impulse:

J=m∆v=F∆t,

m(v_f-v_i )=F∆t,

here, m=100 kg is the mass of the car, v_i=22 m⁄s is the initial velocity of the car, v_f=0 is the final velocity of the car, F is the force on the moving car and ∆t=1.5 s is the time interval during which the car slows down.

Then, from this formula we can find the force on the moving car:

F=m(v_f-v_i )/∆t=(100 kg∙(0 m/s-22 m/s))/(1.5 s)=-1467 N.

The sign minus indicates that the force on the moving car is directed opposite to the motion of the car.

Answer:

F=1467 N, in the opposite direction to the motion of the car.